Two concentric conducting spheres of radii R and 2R are carrying charges Q and -2Q respectively. If the charge on the inner sphere is doubled, the potential difference between the two spheres will be?

Ans. two times

Grade:Upto college level

1 Answers

Akarsh Aithal
10 years ago
since potential difference is directly proportional to the charges then V proportional to Q1*Q2 V1/V2 = Q1*Q2/Q`1*Q`2 V1/V2 = Q*-2Q/2Q*-2Q On simpliyfing V1/V2 = 2/4 V1/V2 = 1/2 2V1=V2